OpenStudy (anonymous):

solve x-19 >(greater then equal to)23 x < (greater then equal to) 42 x >(greater then equal to) 42 x > (greater then equal to) 4 x < (greater then equal to)4

OpenStudy (jdoe0001):

how would you solve x-19 = 23? what would "x" be there?

OpenStudy (anonymous):

4

OpenStudy (jdoe0001):

well, you do the same, just for this specific, pretend the \(\large \ge\) is a =

OpenStudy (anonymous):

so its x> greater then equal to 4

OpenStudy (jdoe0001):

hmm, well, how did you get "4"?

OpenStudy (jdoe0001):

for one, say, x-19 = 23 (+19 both sides) to isolate "x" x= ?

OpenStudy (anonymous):

its 42

OpenStudy (jdoe0001):

:) so \(\arge x\ge 42\)

OpenStudy (anonymous):

how do i know if it goes < or >

OpenStudy (jdoe0001):

woopsie

OpenStudy (anonymous):

*<

OpenStudy (jdoe0001):

so \(\large x\ge 42\)